OpenStudy (mathlegend):

Trig: Angular & Linear Velocity Determine the number of rpm of a wheel turning 52.8 rad/s

OpenStudy (mathlegend):

@ranga can you help me? I have no clue what I am doing when it comes to this stuff.

OpenStudy (mathlegend):

I got the answer of 8.40 but I saw the answer key and it says the answer is: 504.2

OpenStudy (ranga):

52.8 radians = 52.8 / (2pi) revolutions that is in 1 second. In i minute, multiply by 60. 52.8 / (2pi) * 60 revolution per minute or rpm.

OpenStudy (kmeis002):

Well radians is an angular unit which compares arc length to radius. 1 radian is the angle it takes to create an arc length equal to the radius. That being said, there are 2pi radians for 1 revolution (360 degrees) So if we have 1 radian/s and we what revolutions/min, we need to do some conversions. Dimensional analysis will be of help here: \[ \ 52.8 \frac{rad}{s}*\frac{60 \text{seconds}}{1 \text{min}}*\frac{1 \text{ revolution}}{52.8 \text{ radians}}\] Notice the seconds and radians will cancel out, and we will be left with the unit of rev/min

OpenStudy (ranga):

There are (2pi) radians in 1 revolution. Therefore, 1 radian = 1/(2pi) revolutions There are 60 seconds in 1 minute. Therefore, 1 second = 1/60 minutes 52.8 radians / second = 52.8 * (1/2pi) revolutions / 1/60 minutes = 52.8 / (2pi) * 60 revolutions / minute.

OpenStudy (ranga):

= 504.2 rpm

OpenStudy (kmeis002):

you'll notice my error clearly, sorry about that. \[ 52.8 \frac{radians}{second}*\frac{60 \text seconds}{1 \text{ minute}}*\frac{1 \text{ revolution}}{2 \pi \text{ radians}} \]

OpenStudy (mathlegend):

What numbers are you multiplying because I am not getting 504.2...

OpenStudy (ranga):

numerator: 52.8 * 60 denominator: 2(pi)

OpenStudy (mathlegend):

2pi = 360 or 2pi = 6.28

OpenStudy (ranga):

2pi = 6.28.

OpenStudy (kmeis002):

2pi = 6.28

OpenStudy (ranga):

In science and engineering you will come across unit conversions a lot. So it will be good to become familiar with how to do these types of problems. Here we essentially need to convert the units: 52.8 radians / sec = x revolutions / minute what is x? 2pi radians = 1 revolution. so 1 radian = 1/(2pi) revolutions. 60 seconds = 1 minute 1 second = 1/60 minutes. 52.8 radians / sec = 52.8 * 1 radian / 1 second = 52.8 * 1/(2pi) revolutions / 1/60 minutes = (52.8 * 60) / (2pi) revolutions / minute or rpm.

OpenStudy (mathlegend):

So we will always have to ... go from Radians to Revolutions to Minutes

OpenStudy (mathlegend):

4. What is the angular velocity in radians per minute of a notch on a wheel that makes 24 rotations per second about its axis?

OpenStudy (mathlegend):

Is this the same?

OpenStudy (mathlegend):

I have to switch it to revolutions?

OpenStudy (ranga):

pretty much the same procedure as before. 24 rotations / second = x radians / minute. What is x? so we have to replace rotations by radians we have to replace seconds by minutes 1 rotation = 2pi radians 60 seconds = 1 minute 1 second = 1/60 minutes. 24 rotations / second = 24 * (1 rotation / 1 second) = 24 * (2pi radians) / (1/60) minutes = 24 * 2pi * 60 radians / minute

OpenStudy (ranga):

\[\frac{ 24~rotaions }{ second } = 24 *\frac{ 1~rotation }{ 1~second } = 24 * \frac{ 2\pi ~ radians }{ \frac{ 1 }{ 60 }~minute } = 24*2\pi*60~\frac{ radians }{ minute }\]

OpenStudy (ranga):

= 9047.79 radians / minute.

OpenStudy (mathlegend):

Is 1 rotation = 1 revolution?

OpenStudy (ranga):

They use the words interchangeably in these problems. But strictly speaking we say: the earth ROTATES on its own axis once a day and it REVOLVES around the sun once a year.

OpenStudy (ranga):

But in these problems 1 rotation = 1 revolution (the words are synonymous)

OpenStudy (mathlegend):

ok

OpenStudy (ranga):

similar problem you may comes across is to convert miles per hour to feet / second. What is 60 mph in feet/sec? Same procedure. 60 miles / hour = 60 * (1 mile / 1 hour) We need to convert miles to feet in the numerator We need to convert hour to second in the denominator. Look up the units and you can do the conversion the same way.

OpenStudy (ranga):

60 * (1 mile / 1 hour) = x feet / second. What is x? 1 mile = 5280 feet 1 hour = 3600 seconds 60 * (1 mile / 1 hour) = 60 * (5280 feet) / (3600 seconds) = 88 feet / second

OpenStudy (mathlegend):

Is there like websites that practice this stuff?

OpenStudy (ranga):

let me check.

OpenStudy (ranga):

This has some unit conversion examples and exercises: http://www2.franciscan.edu/academic/mathsci/mathscienceintegation/MathScienceIntegation-620.htm

OpenStudy (mathlegend):

Thank You

OpenStudy (ranga):

Here is another: http://www.purplemath.com/modules/units.htm You are welcome.